OpenStudy (anonymous):

Find the solutions of the system: y = 2(5)-x and x + y = 2 0, 2) (0, 2) and approximately (0.09, 1.91) (0, 2) and approximately (1.91, 0.09) no solution

OpenStudy (anonymous):

Fill in the numbers for the original. 0=2(5)-2 <-- no 1.91=2(5)-0.09<-- no So probably no solution. I would get a second opinion though